OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -220xᶾy² -------- 44x¹⁰y⁹

OpenStudy (anonymous):

= -110 / 22x^7y^7 = -55 / 11x^7y^7 = -5/x^7y^7

OpenStudy (anonymous):

thank you